ADNAN MUHAMMED

Grade 12,

A bus started from rest with a new acceleration 2 m/s^2.Show that, a man running with a velocity 10m/s^1 can never catch the bus if he is more than 25 metre behind it.

A bus started from rest with a new acceleration 2 m/s^2.Show that, a man running with a velocity 10m/s^1 can never catch the bus if he is more than 25 metre behind it.

Grade:10

2 Answers

Shivam Shekhar
6 years ago
Let the Separation between them be LVt=1/2at^2+LFrom values....10t+L=t^2For L being max dL/dt should be zero Hence by differentiation 2t=10t=5Hence L max=25
Shivam Shekhar
6 years ago
t=5 (after 2t=10)______________________________________________________________________________________________________________6_6__________________